OpenStudy (anonymous):

The radius of a circular park is 114 yd. To the nearest yard, what is the circumference of the park?

OpenStudy (anonymous):

What will be the diameter if you have the radius?

OpenStudy (anonymous):

I don't know honestly.

OpenStudy (anonymous):

diameter = 2r

OpenStudy (anonymous):

pi = circumference/ diameter

OpenStudy (anonymous):

so 2r = 114yd

OpenStudy (anonymous):

2*114

OpenStudy (anonymous):

228!

OpenStudy (anonymous):

Yes now you have the diameter, can you figure out the circumference?

OpenStudy (anonymous):

3.14?

OpenStudy (anonymous):

That's pi

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

Pi = Circumference / Diameter

OpenStudy (anonymous):

Use it

OpenStudy (anonymous):

I thought the value of a circumference equals 3.14 O__o I don't know what to plug in

OpenStudy (anonymous):

No, ok listen

OpenStudy (anonymous):

pi is the ratio of circumference / diameter ok?

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

So if pi = c/d

OpenStudy (anonymous):

and you're looking for c, I gave you that ratio and that will tell you the answer, I'll give you another hint, ALGEBRA

OpenStudy (anonymous):

you have D now since d = 2*r

OpenStudy (anonymous):

I got 358?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

pi*d = C

OpenStudy (anonymous):

wait hold on..I made a mistake

OpenStudy (anonymous):

Just plug in the numbers now.

OpenStudy (anonymous):

Tell me what you get.

OpenStudy (wolf1728):

Try using this formula: circumference = 2 * PI * radius circumference = 2 * PI * 114 yards

OpenStudy (anonymous):

716?

OpenStudy (anonymous):

:) You got it

OpenStudy (anonymous):

thanks batman ;) & wolf1728!

OpenStudy (wolf1728):

Very good ChinaGurl

OpenStudy (anonymous):

Remember this though, Pi is the ratio of circumference / diameter